# CAPE: A CLIP-Aware Pointing Ensemble of Complementary Heatmap Cues for Embodied Reference Understanding

### v1

@@ Abstract
We address the problem of Embodied Reference Understanding, which involves predicting the object that a person in the scene is referring to through both pointing gesture and language. Accurately identifying the referent requires multimodal understanding: integrating textual instructions, visual pointing, and scene context. However, existing methods often struggle to effectively leverage visual clues for disambiguation. We also observe that, while the referent is often aligned with the head-to-fingertip line, it occasionally aligns more closely with the wrist-to-fingertip line. Therefore, relying on a single line assumption can be overly simplistic and may lead to suboptimal performance. To address this, we propose a dual-model framework, where one model learns from the head-to-fingertip direction and the other from the wrist-to-fingertip direction. We further introduce a Gaussian ray heatmap representation of these lines and use them as input to provide a strong supervisory signal that encourages the model to better attend to pointing cues. To combine the strengths of both models, we present the CLIP-Aware Pointing Ensemble module, which performs a hybrid ensemble based on CLIP features. Additionally, we propose an object center prediction head as an auxiliary task to further enhance referent localization. We validate our approach through extensive experiments and analysis on the benchmark YouRefIt dataset, achieving an improvement of approximately 4 mAP at the 0.25 IoU threshold.
@@ 1 Introduction
While grounding models [37] have made significant progress in detecting objects mentioned in natural language, they often fall short in ERU task, particularly in ambiguous scenes. When multiple instances of the same object type are present, these models tend to detect all matching candidates without the ability to disambiguate and identify the specific target intended by the user. Moreover, when the textual instruction itself is vague or ambiguous, grounding models struggle even further, often failing to identify the correct target object or producing no confident prediction at all due to the inherent ambiguity. This limitation highlights the need for additional disambiguation cues, embodied gesture signals, that can help resolve referential ambiguity and enable accurate identification of the intended object. Therefore, both embodied gesture signals and language reference are crucial to identify the referent. In ERU, two main challenges arise. (1) identifying candidate objects from the text, as in standard grounding tasks, and (2) interpreting the pointing cue from visual input to guide the final prediction. The latter requires understanding human pose, inferring pointing direction, and handling visual complexities like perspective, occlusion, and depth.
In [11], first time ERU task is defined and a multimodal YouRefIt dataset is proposed. They also proposed a method that combines textual and visual inputs along with PAF [7] and predicted saliency maps [33] to help localize the human and narrow down the potential referent. While these inputs lead to improved performance, the model still exhibits suboptimal accuracy in many cases, particularly due to its inability to properly follow the pointing instructions. Building on this, the Touch-Line Transformer [35] introduces a novel approach by predicting a Virtual Touch-Line and applying losses that encourage the model to focus on the head-to-fingertip direction, which they find often aligns closely with the intended referent [24, 44]. Although this method achieves significant gains, it still faces challenges. The model occasionally overlooks the pointing line, as it is not explicitly provided as input. Instead, it predicts head, fingertips, and arm position to encourage attention in that direction. More importantly, the head-to-fingertip direction does not always align with the target. In some scenarios, the accurate pointing direction follows a line from the hand alone (e.g., wrist-to-fingertip), rather than from the head (Fig. 1). These observations suggest that while head-to-fingertip direction offers valuable guidance, it is not sufficient on its own to fully resolve the complexity of the task.
@@
To tackle these challenges, we propose a dual-model framework that leverages complementary pointing cues. Both models take as input a referring expression and an image with a pointing gesture, along with an additional heatmap that encodes the pointing direction. One model utilizes a head-to-fingertip heatmap, while the other uses a wrist-to-fingertip heatmap, capturing different aspects of the pointing behavior. Their predictions are integrated via a CLIP-Aware Pointing Ensemble (CAPE) module, which effectively combines the strengths of both models. Additionally, we introduce an auxiliary objective to predict the object center, providing an extra supervisory signal that further enhances grounding accuracy. Our contributions are: (1) We propose using heatmap as extra modality to guide our end-to-end reference object detection model to focus on pointing information more. (2) We propose two parallel complementary models to focus on head-to-fingertip and wrist-to-fingertip pointing lines, to individually alleviate inherent challenges of the task. (3) We present CAPE module to effectively ensemble two complementary models. (4) We introduce object center prediction head to enhance prediction accuracy by focusing on finding the object without considering the precise bounding box.
@@ 2 Related Work
Some studies [55, 56, 57, 67, 6, 66, 25, 45, 61, 29, 53, 64, 63, 20, 60, 68, 58, 62, 59, 21, 51] use gaze target detection as a nonverbal cue, however, gaze alone is unreliable due to distractions and the absence of a clear pointing moment. It is typically localized frame by frame and often used to support conversation [4, 8]. In [43], a multimodal exophora resolution method is proposed to disambiguate demonstrative expressions like “that one” by integrating object categories, pointing gestures, and prior environmental knowledge. [14] introduced an interactive robot dialogue system that uses multimodal interaction and pointing line estimation to accurately identify referent with an iterative correction process. Their follow-up work [15] extended this by recognizing unseen objects using a Region Proposal Network and VL-T5 multimodal network [13], moving beyond general object classes from benchmarks like MSCOCO [36] to attribute-based object recognition. On the other hand, several datasets have been collected to incorporate nonverbal cues, but most remain in simulated domains due to accessibility and controllability [26, 27, 2, 28].
@@ 3 Methodology
Problem Definition In ERU, given an RGB image and a text input , we predict a bounding box corresponding to an object referenced by the text and indicated by a pointing gesture in the image. We propose using an additional pointing heatmap to supervise the model better focus on the visual pointing reference.
@@ 3.1 Model Architecture
Fig. 2 illustrates our architecture, which consists of a large-scale text encoder, image encoder, and heatmap encoder. The outputs of these encoders are concatenated and fed into a transformer [65] encoder block. Subsequently, a transformer decoder processes the encoder output to generate predictions through task-specific heads. In this work, we introduce two networks that share the same architecture but are trained independently. The first network () receives a heatmap generated from the wrist-to-fingertip line, while the second () uses a heatmap derived from the head-to-fingertip line. Finally, a CLIP-Aware Pointing Ensemble (CAPE) module combines the outputs of these two complementary models to enhance overall performance.
@@
Transformer encoder
After obtaining the embeddings from three encoders, we first apply a convolution to each embedding individually to project their channel dimensions to . Next, we flatten the spatial dimensions of the embedding into a single dimension, converting them into sequences of tokens. For example, for the heatmap embedding: . We then concatenate the three embeddings at the sequence level to form the final representation, which is fed into the transformer encoder. It processes this concatenated input to learn multimodal representations.
Transformer decoder
We feed the multimodal representation output from the transformer encoder into the transformer decoder. Additionally, we provide a set of learnable object queries and gestural keypoint queries. The transformer decoder is responsible for generating object output embeddings and gestural output embeddings.
@@ 3.2 Pointing Heatmap Modality Learning
How should we create heatmap?
The VTL [35] demonstrates superior performance in most cases, as it effectively captures the intended pointing direction by leveraging both the fingertip position and the head orientation. Human pointing gestures are typically guided by visual attention: we often align our finger with the target object based on our eye-level perspective [35]. Consequently, the head–to-fingertip line serves as a strong proxy for the true pointing intent, integrating both gaze and gesture cues. However, the VTL is not universally reliable. In certain scenarios, the performer may be looking elsewhere (for instance, toward a robot or camera) rather than at the intended object. Additionally, in cases involving extreme viewing angles or occluded head poses, the head orientation may become ambiguous or misleading. In such situations, the local orientation of the hand, as approximated by the wrist–to-fingertip line, can offer complementary directional cues that help disambiguate the pointing gesture. We hypothesize that heatmaps derived from the head-to-fingertip and wrist-to-fingertip lines provide complementary information. Therefore, we generate heatmaps that incorporate both information. For the head-to-fingertip heatmap (), we use the eye coordinates as the starting point and draw a line toward the edge of the image, passing through the fingertip. Similarly, for the wrist-to-fingertip heatmap (), we start from the wrist and draw a line to the edge of the image through the fingertip point. To generate these heatmaps, we apply a Gaussian Ray Heatmap (See Appendix B).
@@
The next important question is how to integrate these heatmaps effectively. We explore two different approaches. In the first approach, we merge the two heatmaps () and feed the combined input into a single network to examine whether the model can implicitly learn to utilize both sources of information and determine the most appropriate direction based on the human pose. In the second approach, we train two identical networks independently and ensemble them: one using the head-to-finger heatmap () and the other using the wrist-to-fingertip heatmap (). We empirically find that the second approach yields better performance. As expected, it allows us to train two complementary networks. Notably, the network trained with the head-to-fingertip heatmap () demonstrates significant improvements. In contrast, the first approach results in only subtle enhancement, and the model fails to effectively leverage both sources of information to determine the pointing direction (see Tab. 3 Setup G). Its behavior closely resembles the head-to-fingertip-only scenario. However, due to the presence of additional wrist-to-fingertip line in the merged heatmap, the model experiences contradictory information, ultimately leading to reduced accuracy than head-to-fingertip heatmap-only setup.
Where should we inject these heatmaps?
We empirically find that the most effective way to leverage heatmaps is to use a separate encoder to extract heatmap embeddings, which are then concatenated with the output of the image encoder. It enables us to preserve the pretrained image encoder and allows the dedicated heatmap encoder to extract rich, modality-specific information.
@@ 3.3 CLIP-Aware Pointing Ensemble (CAPE)
We introduce an inference time ensemble module to increase the performance by benefiting from two complementary models, and . CAPE is an adaptive scoring method that combines model confidence and CLIP-based similarity in a size-aware manner effectively. (1) For each prediction (top-2 predictions from both models), we compute a normalized CLIP score [48]. We sum pointing models’ confidece scores with the CLIP scores to obtain the fused score (CLIP Fusion in Tab. 5). (2) We calculate CLIP scores for the top-1 predictions of both models. If the confidence scores of the second-highest predictions exceed a threshold, we also compute their CLIP scores. Finally, we select the prediction with the highest CLIP score from among these candidate boxes (CLIP-Only Top-2 + Threshold=0.95 in Tab. 5). CAPE applies strategy (1) when the referent is a small object (defined as occupying less than 0.48% of the image area, following [11]) to avoid relying solely on the CLIP prediction, as CLIP becomes less reliable for smaller objects. For all other objects, strategy (2) is used (see Sec. C.5 for details). This hybrid approach leverages the strengths of both signals while adapting to object scale. All thresholds and selection rules were tuned on the validation set to avoid test-time overfitting.
@@ 3.4 Gestural Signal Learning
where () and () are the eye and the fingertip coordinates. () are the center of the predicted bounding box. We compute the same similarity using GT bounding box to obtain . The referent alignment loss then penalizes deviations between the predicted and GT alignment:
@@
where is bounding box loss and it is the combination of L1 and GIoU losses as in DETR-like methods [9, 34, 41, 74, 75, 37]. denotes referent alignment loss, indicates object center prediction loss, and is gesture prediction loss. and are soft token loss and contrastive loss respectively to help visual and textural signals alignment as in [30]. We empirically determine the optimal coefficients on the validation set (See Appendix E).
@@ 4 Experimental Results
For both training and testing, we use the image subset of the YouRefIt dataset [11], as the video portion is not publicly available. The dataset contains 2,950 training images and 1,245 test images. Annotations for bounding boxes, as well as eye, fingertip, elbow, and wrist coordinates, are provided by [11, 35].
@@
For fair evaluation and comparison, we follow prior work [11] and report mean Average Precision (mAP) under three Intersection-over-Union (IoU) thresholds: 0.25, 0.50, and 0.75. A prediction is considered correct if its IoU with the GT exceeds the given threshold. Additionally, by following the literature, mAP scores are reported with respect to object size, categorized as Small (S), Medium (M), and Large (L). The object size categories are determined based on the ratio of the object area to the image area, using two thresholds [11]: and . An object is classified as small if its area ratio is below , medium if it falls between and , and large otherwise. In addition to this standard evaluation protocol, we introduce two metrics for assessing referential grounding performance: CLIP score and center coordinate distance (). To compute the CLIP score, we extract features from the text input and the cropped region corresponding to the predicted bounding box using CLIP’s text and image encoders. We then calculate the cosine similarity between these feature vectors to quantify the semantic alignment between the predicted object and the text. For , we compute the L1 distance between the center coordinates of the predicted bounding box and those of the GT bounding box, providing a direct measure of spatial alignment.
@@
| Touch-Line-EWL | 0.7369 | 0.6936 | 0.7305 | 0.7845 | 0.3168 | 0.3006 | 0.2903 | 0.3564 |
| Touch-Line-VTL | 0.7368 | 0.6925 | 0.7321 | 0.7846 | 0.2843 | 0.2809 | 0.2276 | 0.3393 |
| Ours | 0.7983 | 0.7927 | 0.8015 | 0.8010 | 0.2476 | 0.2137 | 0.2241 | 0.3023 |

### v4

@@ Abstract
We address Embodied Reference Understanding, the task of predicting the object a person in the scene refers to through pointing gesture and language. This requires multimodal reasoning over text, visual pointing cues, and scene context, yet existing methods often fail to fully exploit visual disambiguation signals. We also observe that while the referent often aligns with the head-to-fingertip direction, in many cases it aligns more closely with the wrist-to-fingertip direction, making a single-line assumption overly limiting. To address this, we propose a dual-model framework, where one model learns from the head-to-fingertip direction and the other from the wrist-to-fingertip direction. We introduce a Gaussian ray heatmap representation of these lines and use them as input to provide a strong supervisory signal that encourages the model to better attend to pointing cues. To fuse their complementary strengths, we present the CLIP-Aware Pointing Ensemble module, which performs a hybrid ensemble guided by CLIP features. We further incorporate an auxiliary object center prediction head to enhance referent localization. We validate our approach on YouRefIt, achieving 75.0 mAP at 0.25 IoU, alongside state-of-the-art CLIP and scores, and demonstrate its generality on unseen CAESAR and ISL Pointing, showing robust performance across benchmarks.
@@ 1 Introduction
While grounding models [39] and LMMs [3, 4, 6, 56] have made significant progress in detecting objects mentioned in natural language, they often fall short in ERU task, particularly in ambiguous scenes. When multiple instances of the same object type are present, these models tend to detect all matching candidates without the ability to disambiguate and identify the specific target intended by the user. Moreover, when the textual instruction itself is vague or ambiguous, grounding models struggle even further, often failing to identify the correct target object or producing no confident prediction at all due to the inherent ambiguity. This limitation highlights the need for additional disambiguation cues that can help resolve referential ambiguity and enable accurate identification of the intended object. Therefore, both embodied gesture signals (pointing) and language reference are crucial to identify the referent, aligning with early multimodal interaction studies [72, 61].
The ERU task was first defined in [13] with the YouRefIt dataset, and a method combining textual and visual inputs with PAF [8] and predicted saliency maps [36] was proposed. While effective, the model often fails to follow pointing instructions accurately. The Touch-Line Transformer [38] addresses this by predicting a Virtual Touch-Line along the head-to-fingertip direction. However, it still faces limitations as the pointing line is not explicitly provided as input, despite its success [84], and the head-to-fingertip direction does not always align with the target. In some cases, accurate pointing is better represented from the hand alone (e.g., wrist-to-fingertip, Fig. 1), highlighting the need for more reliable pointing guidance.
@@
To tackle these challenges, we propose a dual-model framework that leverages complementary pointing cues. Both models take as input a referring expression and an image with a pointing gesture, along with an additional heatmap that encodes the pointing direction. One model utilizes a head-to-fingertip heatmap, while the other uses a wrist-to-fingertip heatmap, capturing different aspects of the pointing behavior. Since these two models provide complementary information depending on the scenario, we introduce the CLIP-Aware Pointing Ensemble (CAPE) module to effectively combine the strengths of both models. In this module, we leverage the CLIP model to compute semantic similarity scores between the input text and candidate object images. We use CLIP because it was trained on a large-scale image-text dataset, learning a joint embedding space in which semantically matching images and text are closely aligned. Additionally, we introduce an auxiliary object center prediction head, providing a supervisory signal that guides the model to more accurately localize the pointed object.
Our contributions are: (1) We propose using a pointing heatmap as an additional modality to guide our model to focus more effectively on pointing cues. Specifically, we use detected head, wrist, and fingertip points to construct a Gaussian Ray Heatmap that highlights the approximate pointing area. The embedding of this pointing heatmap is then extracted via a heatmap encoder and provided to the model. (2) We propose two parallel complementary models to focus on head-to-fingertip and wrist-to-fingertip pointing lines, to individually alleviate inherent challenges of the task. (3) We present CAPE module to effectively ensemble two complementary models as ensembling has shown strong performance in other fields [19, 82, 27, 78, 83]. (4) We introduce an object center prediction head to improve prediction accuracy by emphasizing object localization without requiring precise bounding boxes.
@@ 2 Related Work
Some studies [57, 58, 59, 60, 71, 7, 70, 28, 46, 65, 32, 54, 68, 67, 23, 64, 73, 62, 66, 63, 24, 52] use gaze target detection as a nonverbal cue, however, gaze alone is unreliable due to distractions and the absence of a clear pointing moment. It is typically localized frame by frame and often used to support conversation [5, 10]. In [45], a multimodal exophora resolution method is proposed to disambiguate demonstrative expressions like “that one” by integrating object categories, pointing gestures, and prior environmental knowledge. [16] introduced an interactive robot dialogue system that uses multimodal interaction and pointing line estimation to accurately identify referent with an iterative correction process. Their follow-up work [17] extended this by recognizing unseen objects using a Region Proposal Network and VL-T5 multimodal network [15], moving beyond general object classes. On the other hand, several datasets have been collected to incorporate nonverbal cues, but most remain in simulated domains due to accessibility and controllability [29, 30, 2, 31].
@@ 3 Methodology
In ERU, given an RGB image and a text input , we predict a bounding box corresponding to an object referenced by the text and indicated by a pointing gesture in the image. We propose using an additional pointing heatmap to supervise the model better focus on the visual pointing reference.
@@ 3.1 Model Architecture
Fig. 2 illustrates our overall approach, which consists of two parallel models ( and ; only one is shown since the sole difference is the heatmap encoder input). After processing the text, image, and heatmap inputs, the outputs of both models are passed to the CLIP-Aware Pointing Ensemble (CAPE) module for the final decision. CAPE combines the outputs of these two complementary models to enhance overall performance. Specifically, and each comprise a large-scale text encoder, image encoder, and heatmap encoder. The outputs of these encoders are concatenated and fed into a transformer [69] encoder block. A subsequent transformer decoder then processes the encoder output to generate predictions via task-specific heads.
@@
Transformer encoder and decoder
After obtaining the embeddings from three encoders, we first apply a convolution to each embedding individually to project their channel dimensions to . Next, we flatten the spatial dimensions of the embedding into a single dimension, converting them into sequences of tokens. For example, for the heatmap embedding: , where is flatten and is convolution. We then concatenate the three embeddings at the sequence level to form the final representation, which is fed into the transformer encoder. It processes this concatenated input to learn multimodal representations. We feed the multimodal representation output from the transformer encoder into the transformer decoder. Additionally, we provide a set of learnable object queries and gestural keypoint queries. The transformer decoder is responsible for generating object output embeddings and gestural output embeddings.
@@ 3.2 Pointing Heatmap Modality Learning
How should we create heatmaps?
While prior work such as VTL [38] shows that head–to–fingertip cues can often capture the intended pointing direction, our goal is to design a heatmap representation that remains reliable across a wider range of real-world conditions. Human pointing gestures are guided by visual attention, and the head–to–fingertip line naturally encodes this gaze–gesture alignment. However, relying on this cue alone can be problematic when the performer is looking elsewhere (e.g., toward a robot or camera) or when the head pose becomes ambiguous due to extreme angles or occlusions. Moreover, close proximity to the object or incomplete arm extension can also break alignment. To address these limitations, our approach incorporates both the head–to–fingertip and wrist–to–fingertip lines. The wrist–to–fingertip line provides a complementary, locally grounded cue that remains informative even when head orientation is unreliable. By combining these two directional sources, we aim to obtain a more robust and consistent representation of pointing intent. Based on this hypothesis, we generate Gaussian Ray Heatmaps for both cues rather than relying solely on explicit line overlap. Both heatmaps extend a ray toward the image boundary through the fingertip, originating at the eyes for and at the wrist for . These heatmaps serve as complementary spatial supervision signals generated through our heatmap process (see Appendix A), capturing not only overlap with a single pointing line but also a broader region that reflects the performer’s focus of attention.
@@
We investigate two strategies for integrating the two heatmaps. The first merges them directly () and feeds the combined input into a single network, testing whether the model can implicitly resolve the two cues and infer the correct pointing direction from pose context. The second trains two identical networks separately, one with the head-to-fingertip heatmap () and one with the wrist-to-fingertip heatmap (), and ensembles their predictions. Empirically, the second strategy works substantially better. It yields two complementary models, with providing the strongest gains. In contrast, the merged-heatmap approach offers only marginal improvement. The model behaves similarly to the head-only variant, and the extra wrist-to-fingertip signal introduces conflicting directional cues that ultimately reduce accuracy compared to using the head-to-fingertip heatmap alone (see Tab. 4, Setup G). We find that the most effective way to leverage heatmaps is to process them with a dedicated encoder whose embeddings are concatenated with the image encoder’s output. This design preserves the pretrained visual backbone while allowing the heatmap encoder to capture richer, modality-specific spatial cues.
@@ 3.3 CLIP-Aware Pointing Ensemble (CAPE)
We introduce an inference time ensemble module to increase the performance by benefiting from two complementary models, and , since we find that they demonstrate strength in different scenarios. CAPE is an adaptive scoring method that combines model confidence and CLIP-based similarity in a size-aware manner effectively. We choose the CLIP model because it was trained to understand and evaluate image-text semantic similarity, which aligns well with our task. (1) For each prediction (top-2 predictions from both models), we compute a normalized CLIP score [49]. We sum pointing models’ confidence scores with the CLIP scores to obtain the fused score (CLIP Fusion in Tab. 6) since the models’ confidence scores also provide reliable insights. (2) We calculate CLIP scores for the top-1 predictions of both models. If the confidence scores of the second-highest predictions exceed a threshold, we also compute their CLIP scores. Finally, we select the prediction with the highest CLIP score from among these candidate boxes (CLIP-Only Top-2 + Threshold=0.95 in Tab. 6). CAPE applies strategy (1) when the referent is a small object (defined as occupying less than 0.48% of the image area, following [13]) to avoid relying solely on the CLIP prediction, as CLIP becomes less reliable for smaller objects. For all other objects, strategy (2) is used (see Sec. B.5 for details). This hybrid approach leverages the strengths of both signals while adapting to object scale. All thresholds and selection rules were tuned on the validation set to avoid test-time overfitting.
@@ 3.4 Gestural Signal Learning
where () and () are the eye and the fingertip coordinates. () are the center of the predicted bounding box. is cosine similarity. We compute the same similarity using GT bounding box to obtain . The referent alignment loss then penalizes deviations between the predicted and GT alignment:
@@
where is bounding box loss and it is the combination of L1 and GIoU losses as in DETR-like methods [11, 37, 43, 81, 85, 39]. denotes referent alignment loss, indicates object center prediction loss, and is gesture prediction loss. and are soft token loss and contrastive loss respectively to help visual and textural signals alignment as in [33]. We empirically determine the optimal coefficients on the validation set (See Appendix D).
@@ 4 Experimental Results
For both training and testing, we use the YouRefIt dataset [13], which contains 2,950 training and 1,245 test images. Text instructions and the annotations of bounding boxes, as well as pointing line coordinates, are provided by [13, 38]. Additionally, we evaluate our approach on the unseen and more challenging ISL Pointing dataset [16], and a simulation dataset named CAESAR [29] (for results see appx. Tab. E.13) to assess its generalization capability.
@@
For fair evaluation and comparison, we follow prior work [13] and report mean Average Precision (mAP) under three Intersection-over-Union (IoU) thresholds: 0.25, 0.50, and 0.75. Additionally, mAP scores are reported with respect to object size, categorized as Small (S), Medium (M), and Large (L) which based on the ratio of the object area to the image area using two thresholds: and . In addition to this standard evaluation protocol, we introduce two metrics for assessing referential grounding performance: CLIP score and center coordinate distance (). To compute the CLIP score, we extract features from the text input and the cropped region corresponding to the predicted bounding box using CLIP’s text and image encoders. We then calculate the cosine similarity between these feature vectors to quantify the semantic alignment between the predicted object and the text. For , we compute the L1 distance between the center coordinates of the predicted and the GT bounding boxes, providing a direct measure of spatial alignment.
@@
| PaliGemma2 | 0.2488 | 0.2302 | 0.2446 | 0.2659 | 0.4684 | 0.5784 | 0.4307 | 0.4175 |
| Qwen2.5vl-3b | 0.2408 | 0.2261 | 0.2388 | 0.2568 | 0.7288 | 0.7074 | 0.6893 | 0.7891 |
| Grounding DINO | 0.2437 | 0.2267 | 0.2413 | 0.2621 | 0.3496 | 0.2946 | 0.4521 | 0.3104 |
| Touch-Line-EWL | 0.2456 | 0.2312 | 0.2435 | 0.2615 | 0.3168 | 0.3006 | 0.2903 | 0.3564 |
| Touch-Line-VTL | 0.2456 | 0.2308 | 0.2440 | 0.2615 | 0.2843 | 0.2809 | 0.2276 | 0.3393 |
| Ours (CAPE) | 0.2661 | 0.2642 | 0.2671 | 0.2670 | 0.2476 | 0.2137 | 0.2241 | 0.3023 |
| Method | IoU=0.25 | IoU=0.5 | IoU=0.75 | CLIP | |
| PaliGemma2 | 47.2 | 39.5 | 31.6 | 0.2465 | 0.7449 |
| Qwen2.5vl-3b | 32.5 | 32.1 | 29.2 | 0.2264 | 0.8418 |
| Grounding DINO | 27.5 | 27.5 | 26.3 | 0.2082 | 0.7956 |
| Touch-Line-EWL | 45.0 | 35.8 | 22.0 | 0.2436 | 0.5160 |
| Touch-Line-VTL | 47.7 | 36.7 | 17.4 | 0.2473 | 0.5147 |
| Ours, | 48.2 | 33.6 | 20.0 | 0.2453 | 0.4735 |
| Ours, | 42.7 | 30.0 | 17.3 | 0.2449 | 0.6795 |
| Ours, CAPE | 54.5 | 42.7 | 25.5 | 0.2642 | 0.4403 |